Consider the same air mass from question 16: the velocity of uplift is measured at 240 meters per
hour. At what time can we expect cloud formation? Be precise in your answer.
1
Expert's answer
2019-10-25T11:22:59-0400

the rate of change of the saturation ratio S is given by:

"\frac{ dS}{dt}=P-S=Q_1×\frac{ dz}{dt} - Q_2×\frac{ dλ}{dt}"


where P= production, C= condensation


"\frac{ dz}{dt}" is the vertical air velocity


"\frac{ dλ}{dt}" is the condensation rate


Time is equal to 2.5 hour